OpenStudy (anonymous):

a/b+c/d=b/a+d/c,then prove that a3/b3+c3/d3=b3/a3+d3+c3.

OpenStudy (anonymous):

help me yeer

OpenStudy (anonymous):

Do you mean d3/c3 not d3+c3?

OpenStudy (anonymous):

wait. i'm doing it for you

OpenStudy (anonymous):

hey dumbcow, can you help me with my question above?

OpenStudy (dumbcow):

For first statement, combine fractions on both sides and simplify leaving bd = ac thus b^3d^3 = a^3c^3 b^3/a^3 = c^3/d^3 substitute into second statement to complete proof

OpenStudy (anonymous):

Now see.

OpenStudy (anonymous):

are you satisfied now?